# Low-Frequency Noise and Resistive Switching in $\beta$-Na0.33V2O5

### v1

@@ I Introduction
Low-dimensional systems with strong electronic correlations exhibit a variety of unconventional phenomena, including charge ordering, polaronic behavior, and metal-insulator transitions.[1, 2, 3] Electronic phase transitions have significant effects on transport properties and are sensitive to subtle modulation of external parameters such as temperature, pressure, or electric field.[4, 5, 6] The competition between localization, electronic correlations, and external driving forces needs to be clarified to disentangle the mechanisms that govern conduction in low-dimensional systems.[6, 7] In the past decade, correlated oxides and mixed-valence compounds have attracted wide attention due to their promise to exhibit tunable resistive states and electric field-driven switching behavior, [8, 9, 10, 11, 12] making them as attractive candidates for fundamental studies of non-equilibrium transport and as next-generation electronic devices that operate across a range of ambient environments.[13, 14, 15, 16, 17] However, a clear understanding of how microscopic electronic phenomena, such as charge ordering, Mott transitions, and polaron formation, influence macroscopic transport properties and resistive switching behavior remains elusive. This gap is especially pronounced in quasi-one-dimensional correlated systems, where the interplay of charge, lattice, and spin disorder leads to complex and emergent electronic responses.[18]
@@
The quasi-one-dimensional $\beta$ ($\beta^{\prime}$)-AxV2O5 (A = Li, Na, Ca, Pb, Cu, etc.) vanadium oxide bronzes represent a class of materials exhibiting resistance changes spanning several orders of magnitude in response to temperature and electric field.[19, 20, 21, 22] In these bronzes, the electrical resistance can be tuned by the choice of intercalant, giving rise to a wide range of electrical and magnetic behaviors within this material family.[23] The crystals feature a distinctive layered tunnel structure, characterized by high anisotropy, enabling novel phenomena such as a high-temperature metallic state observed when specific stoichiometric ratios of inserted cations, typically $x=0.33$ or $0.65$ for Cu insertion in the $\beta^{\prime}$-phase, are incorporated into the tunnel.[24] For Na inserton at $x=0.33$, the average structure contains one Na-ion randomly distributed within every pair of neighboring $\beta$-sites, resulting in a half-occupied Na-ion sublattice stabilized by local zig-zag ordering.[25] This arrangement induces a long-range periodic potential parallel to the $b$-axis throughout the tunnel structure, giving rise to a quasi-one-dimensional metallic state at room temperature.[23, 24] However, this metallic state is highly sensitive to off-stoichiometry and crystal defects, which disrupt the Na-induced long-range potential and pattern of electron and spin localization along the host $\zeta$-V2O5 framework, which significantly affects the material’s electrical properties.
@@
Additionally, these vanadium bronzes are mixed-valence compounds, with V4+ (3d1) and V5+ (3d0) distributed between the V1 and V2 sites. These sites exhibit octahedral coordination with the adjacent oxygen atoms. The (V1)O6 octahedra form a zig-zag double-chain structure, whereas the (V2)O6 octahedra create a two-legged ladder structure, aligned parallel to the $b$-axis. The non-magnetic V3 site, which has a formal valence of V5+, forms zig-zagging square pyramidal coordination, also parallel to the $b$-axis. This unique arrangement leads to the tunnel structure observed in these materials, where the $\beta$-inserted-ion resides within the $\beta$-sites in the tunnels, as illustrated in Fig. 1.[23, 26, 27, 24, 28, 29] Notably, the $\beta$-sites cannot have both nearest-neighbor positions occupied simultaneously, because of cationic repulsions. The $\beta^{\prime}$-sites, which are always farther apart than the $\beta$-sites, are typically occupied when cation stoichiometry exceeds 0.33 as in the case of Li- and Cu-inserted $\beta^{\prime}$-AxV2O5 compounds.
@@
To explore correlated charge phenomena, low-frequency resistance noise spectroscopy has proven to be a powerful probe for detecting dynamic fluctuations and phase transitions in complex materials.[18, 30, 31] Unlike conventional transport measurements, resistance noise reveals hidden information about carrier dynamics, disorder, and correlation effects. In particular, a sharp drop in noise magnitude is recognized as a signature of collective or glassy charge dynamics in correlated systems.[32] In parallel, charge ordering in low-dimensional oxides has been linked to field-induced resistive switching, highlighting the role of electronic phase transitions in enabling non-volatile conductance states.[33] These insights motivate a deeper investigation of switching mechanisms in charge-ordered materials such as $\beta$-Na0.33V2O5.
@@
Upon cooling from room temperature, $\beta$-Na0.33V2O5 undergoes a sequence of phase transitions. Initially, the Na-ions, which occupy one of two distinct crystallographic positions within the tunnel network, [26] begin to structurally order at $T_{\text{IO}}\sim 230$ K, [34, 26] forming a zig-zag pattern that alternates between $\beta$-sites along the $b$-axis. As the temperature is further reduced below $T_{\text{IO}}$, the Na-ion ordering becomes complete, eventually leading to charge redistribution primarily onto the V1 zig-zag chains.[29, 35] Upon further cooling, a charge ordering (CO) transition occurs at $T_{\text{CO}}\sim 130$ K, [28, 34] which is accompanied by the localization of carriers. Optical spectroscopy suggests that these charge carriers may form small polarons, underscoring the significant role of electron–phonon interactions in stabilizing the CO ground state.[36] At $T_{\text{N}}\sim 22$ K, the system undergoes spin ordering, forming a canted antiferromagnetic structure.[37] Finally, under hydrostatic pressures exceeding 8 GPa, a superconducting phase emerges below 10 K.[24, 37, 27, 38] The coexistence of strong correlations, reduced dimensionality, and tunable ground states render this system a fertile ground for fundamental studies and for exploring its potential as a platform for memory and computing applications, motivating further investigation of its complex phase behavior.
@@ II Experimental Methods
Single crystals of $\beta$-Na0.33V2O5 were prepared by a solid-state powder synthesis and subsequent melt-growth, as described in .[39] Na2C2O4 (Sigma-Aldrich) and V2O5 (Sigma-Aldrich) at a 0.18:1 molar ratio (a small stoichiometric excess of Na was used to maximize Na-site filling) were ground together thoroughly in a mortar and pestle and loaded into an alumina combustion boat. The mixture was placed in an ambient-temperature tube furnace, held under vacuum, and heated to 550°C for 16 h. After cooling, the sample was removed, ground, and heated as previously. The identity of the resulting blue-black powder was confirmed by powder X-ray diffraction. To prepare single crystals, the powder product was sealed in an evacuated amorphous silica ampoule, melted at 810°C in a muffle furnace, cooled to 550°C at 2°C/h to crystallize, and cooled to ambient temperature under furnace momentum. Several regular, lustrous black acicular crystals were harvested for measurement.
@@ III RESULTS AND DISCUSSION
Above $T_{CO}$, the data were fit to the nearest-neighbor hopping (NNH) model for small polarons by plotting $\ln(\rho/T)$ versus $T^{-1}$, as shown in Fig. 2(b), following Eq. 1.[35] The extracted activation energy is approximately 63 meV, in agreement with previously reported values for $\beta$-Na0.33V2O5.[42, 38] This energy represents the thermal barrier for polaron hopping and reflects the degree of carrier localization, making it a critical parameter for understanding the transport mechanism in the high-temperature regime.
@@
However, as shown in Fig. 2(b), the NNH fit is valid only in the temperature range from 130 K to 220 K, suggesting the presence of another transition near this range, which is discussed later. The variable-range hopping (VRH) conduction mechanism below $T_{CO}$ is analyzed by plotting $\ln(\rho/T^{1/2})$ against $T^{-1/4}$, following the temperature dependence described by Eq. 2, [42, 43] as depicted in Fig. 2(c).
@@
This analysis yields Mott’s characteristic temperature $T_{0}$, which is found to be $1.88\times 10^{8}\,\text{K}$. Based on this value, the density of states at the Fermi level $N(E_{f})$, the mean hopping distance $R$, and the mean hopping energy $W$ were calculated. Assuming that the small polaron is localized to a single vanadium site, the localization length was approximated as $\alpha^{-1}\sim 1\,\text{Å}$.[43] These calculations yielded $N(E_{f})=1.16\times 10^{21}\,\text{eV}^{-1}\text{cm}^{-3}$, $R=1.27\,\text{nm}$, and $W=100.47\,\text{meV}$. Additionally, for the VRH model to remain valid, the conditions $\alpha R>1$ and $W>k_{B}T$ were verified for all temperatures below $T_{CO}$, as outlined in Eq. 3.[42, 44]
@@
To probe the carrier dynamics across the transition, low-frequency noise measurements were performed over the temperature range of 100 K to 300 K. Fig. 2(d) displays representative residual current fluctuations measured above, near, and below the transition, with notably higher fluctuations observed in the vicinity of $T_{\text{CO}}$. The PSD of these fluctuations are shown in Fig. 2(e), exhibiting a clear inverse frequency dependence. Fig. 2(a) shows the PSD magnitude at 0.1 Hz as a function of T, revealing a weak temperature dependence above $T_{\text{CO}}$ and a sharp suppression of noise as the system enters the charge-ordered state below $T_{\text{CO}}$. This weak temperature dependence at high-temperatures is characteristic of hopping conduction in systems [45, 46] where transport is dominated by thermally activated hopping between nearest-neighbor localized states. In this regime, the resistivity becomes nearly temperature-independent, and the conduction pathways stabilize, leading to minimal variation in noise magnitude. The strong suppression of noise below $T_{\text{CO}}$, in contrast, reflects the onset of enhanced carrier localization and collective electronic correlations-providing noise-based evidence for the transport crossover from NNH to VRH, as suggested by the resistivity analysis.
Similar to optical measurements showing a gradual ordering of the Na-ions, [29] further insights into the temperature dependence of the IO and CO can be probed through X-ray diffraction intensities, as shown in Fig. 3 (c). These intensities reveal a gradual increase in CO due to enhanced electron localization as the system is further cooled below $T_{CO}$. This progressive charge localization at lower temperatures underpins for the observed pressure-temperature (PTP-T) diagrams in many other organic and heavy-fermion systems exhibiting pressure-induced superconductivity.
Recent reports have also discussed an additional transition corresponding to Na-ion ordering (IO) at $T_{IO}\sim 240\,\text{K}$.[47, 48] Optical measurements have shown that this transition occurs gradually from $T_{IO}$ to $T_{CO}$, where charge localization occurs along the zig-zag chains of vanadium atoms spanning the $b$-axis, centered on the $V_{1}$ vanadium atom.[34] Single-crystal X-ray diffraction measurements reveal the atomic structure modulations exhibited by the low-temperature insulating state. $\beta$-Na0.33V2O5 exhibits the C2/m space group at room temperature, shown in Fig. 3 (c). (V1)O6 and (V2)O6 octahedra form V4O16 clusters, which edge-share in chains along $b$, connected along $a$ by shared V1-O2-V1 corners and bridged along $c$ by zig-zag chains of edge-shared (V3)O5 square pyramids. Na-ions occupy paired sites within tunnel-shaped interstices extending along $b$. The proximity of these site pairs (1.953 Å) limits them to half-occupancy, corresponding to a nominal Na1/3V2O5 composition.
@@
Below $T_{\text{CO}}=125$ K, the material adopts an ($a$, $6b$, $c$) supercell, as indicated by the presence of $\mathbf{q}=(h,k/6,l)$ satellite reflections at 100 K in Fig. 3 (b). Vanadium bond valence sum [49] calculations suggest the formation of an approximate $3b$ periodic CO in conjunction with the extant $2b$ Na ordering as the source of the modulation ( Fig. 3 (a) ). CO has previously been invoked to describe the conductivity and structural transitions and others coincident at 125 K, [28, 29] and several charge distribution models have been proposed.[38, 50, 27] Notably, the charge distribution present is nearly identical to that supporting the antiferromagnetic ground state, as determined by single-crystal neutron diffraction below the Néel temperature.[51] This magnetic ordering likely coalesces from the CO pattern established at the onset of the semiconductor-insulator transition as presented above. Like the lower-order satellites, Fig. 3 (c) shows that the reflections at $\mathbf{q}=(h,k/6,l)$ intensify with cooling, indicating deepening of charge localization.
@@
In non-interacting systems, noise increases as the temperature decreases due to the narrowing of the density of states near the Fermi energy ($E_{f}$) at lower temperatures, which elevates the energy barriers for hopping. As a result, hopping events between localized states require longer timescales, leading to higher noise magnitudes. Essentially, the reduced availability of states for particles to hop causes higher noise levels at lower temperatures, characteristic of non-interacting systems.[46, 52] In contrast to characteristic non-interacting patterns, the observed noise behavior in single crystals of $\beta$-NaxV2O5 exhibits pronounced differences. Indeed, the noise magnitude decreases below the characteristic transition temperature TCO, a phenomenon that cannot be explained within the framework of non-interacting models. The observed anomaly is attributed to the onset of strong electron-electron correlations due to CO, which drive a collective reorganization of localized states under the influence of Coulomb interactions. In this correlated regime, these interactions suppress the independent hopping fluctuations characteristic of non-interacting systems, ultimately giving rise to a glassy state.[53, 54] As previously reported in optical studies, the observed mid-infrared absorption, combined with strong electron-phonon coupling and effective carrier density, suggests that small polaron transport plays a dominant role in $\beta$-Na0.33V2O5.[29] Hall measurements have revealed that the total number of charge carriers remains largely unaffected by temperature changes in the range of 189 K to 298 K, with the carrier density approximately matching the concentration of Na atoms in the crystal.[55] This gives additional support to the polaronic nature of conduction, as each excess electron introduced into the lattice by the intercalated monovalent Na localizes on the $V_{1}/V_{2}$ vanadium sites, which exhibit a mixed valence of +4/+5.
@@
Further analysis is presented in Fig. 5(c) and (d). In Fig. 5(c), the noise magnitude at $f=10~\mathrm{mHz}$ remains low at low voltages, but increases sharply as the system approaches $V\sim 16.5~\mathrm{V}$, peaking just before the current discontinuity. The abrupt rise abrupt rise in noise is indicative of critical dynamics often observed during the collapse of an ordered electronic state. In the context of $\beta$-Na0.33V2O5, such dynamics are consistent with the field-driven destabilization of the charge-ordered state formed by localized small polarons. It suggests a regime where the applied voltage competes with the intrinsic pinning potential imposed by lattice or impurity interactions. Simultaneously, Fig. 5(d) shows the evolution of the PSD exponent $\alpha$, which is close to 1 at low voltages-typical of $1/f$ noise. As the transition is approached, $\alpha$ increases toward 2, reflecting a shift to noise behavior dominated by slow, diffusive-like dynamics. A $1/f^{2}$ noise spectrum ($\alpha\approx 2$) is indicative of non-equilibrium dynamics associated with depinning or cooperative motion. Similar behavior has been observed in charge-ordered organic conductors such as $\kappa$-(BETS)2Mn[N(CN)2]3, where a sharp rise in noise and $\alpha\approx 2$ appears below the transition, reflecting slow, correlated fluctuations of polar nanoregions.[56]
@@
Moreover, the temperature dependence of the threshold voltage $V_{\text{th}}$ shown in Fig. 6(b), displays an exponential behavior. This trend is consistent with prior observations in related charge-ordered systems,[57, 58] where the CO phase becomes increasingly susceptible to electric-field-induced breakdown at lower temperatures.

### v2

@@ I Introduction
Low-dimensional systems with strong electronic correlations exhibit a variety of unconventional phenomena, including charge ordering, polaronic behavior, and metal-insulator transitions.[alder2018, Moser2013, sood2021] Electronic phase transitions have significant effects on transport properties and are sensitive to subtle modulation of external parameters such as temperature, pressure, or electric field.[Lee2018, vale2019, dagotto2005] The competition between localization, electronic correlations, and external driving forces needs to be clarified to disentangle the mechanisms that govern conduction in low-dimensional systems.[dagotto2005, Alam2023] In the past decade, correlated oxides and mixed-valence compounds have attracted wide attention due to their promise to exhibit tunable resistive states and electric field-driven switching behavior, [Nitin2025, sawa2008, waser2009, janod2015, kaushik2025] making them as attractive candidates for fundamental studies of non-equilibrium transport and as next-generation electronic devices that operate across a range of ambient environments.[Hutter_2011, Jo2010, lanyon2008, Pathak2018, Preskill_2018] However, a clear understanding of how microscopic electronic phenomena, such as charge ordering, Mott transitions, and polaron formation, influence macroscopic transport properties and resistive switching behavior remains elusive. This gap is especially pronounced in quasi-one-dimensional correlated systems, where the interplay of charge, lattice, and spin disorder leads to complex and emergent electronic responses.[Geremew2019_ACSNano]
@@
The quasi-one-dimensional $\beta$ ($\beta^{\prime}$)-AxV2O5 (A = Li, Na, Ca, Pb, Cu, etc.) vanadium oxide bronzes represent a class of materials exhibiting resistance changes spanning several orders of magnitude in response to temperature and electric field.[Marley2014, PARIJA20201166, Marley2013, Andrew2019] In these bronzes, the electrical resistance can be tuned by the choice of intercalant, giving rise to a wide range of electrical and magnetic behaviors within this material family.[Ma2006] The crystals feature a distinctive layered tunnel structure, characterized by high anisotropy, enabling novel phenomena such as a high-temperature metallic state observed when specific stoichiometric ratios of inserted cations, typically $x=0.33$ or $0.65$ for Cu insertion in the $\beta^{\prime}$-phase, are incorporated into the tunnel.[Okazaki2004] For Na inserton at $x=0.33$, the average structure contains one Na-ion randomly distributed within every pair of neighboring $\beta$-sites, resulting in a half-occupied Na-ion sublattice stabilized by local zig-zag ordering.[Krogstad2020] This arrangement induces a long-range periodic potential parallel to the $b$-axis throughout the tunnel structure, giving rise to a quasi-one-dimensional metallic state at room temperature.[Ma2006, Okazaki2004] However, this metallic state is highly sensitive to off-stoichiometry and crystal defects, which disrupt the Na-induced long-range potential and pattern of electron and spin localization along the host $\zeta$-V2O5 framework, which significantly affects the material’s electrical properties.
@@
Additionally, these vanadium bronzes are mixed-valence compounds, with V4+ (3d1) and V5+ (3d0) distributed between the V1 and V2 sites. These sites exhibit octahedral coordination with the adjacent oxygen atoms. The (V1)O6 octahedra form a zig-zag double-chain structure, whereas the (V2)O6 octahedra create a two-legged ladder structure, aligned parallel to the $b$-axis. The non-magnetic V3 site, which has a formal valence of V5+, forms zig-zagging square pyramidal coordination, also parallel to the $b$-axis. This unique arrangement leads to the tunnel structure observed in these materials, where the $\beta$-inserted-ion resides within the $\beta$-sites in the tunnels, as illustrated in Fig. 1.[Ma2006, Markina2004, Itoh2006, Okazaki2004, van2002, Presura2003] Notably, the $\beta$-sites cannot have both nearest-neighbor positions occupied simultaneously, because of cationic repulsions. The $\beta^{\prime}$-sites, which are always farther apart than the $\beta$-sites, are typically occupied when cation stoichiometry exceeds 0.33 as in the case of Li- and Cu-inserted $\beta^{\prime}$-AxV2O5 compounds.
@@
To explore correlated charge phenomena, low-frequency resistance noise spectroscopy has proven to be a powerful probe for detecting dynamic fluctuations and phase transitions in complex materials.[Geremew2019_ACSNano, singh2016, Alsaqqa2017] Unlike conventional transport measurements, resistance noise reveals hidden information about carrier dynamics, disorder, and correlation effects. In particular, a sharp drop in noise magnitude is recognized as a signature of collective or glassy charge dynamics in correlated systems.[Samanta2012_PRB] In parallel, charge ordering in low-dimensional oxides has been linked to field-induced resistive switching, highlighting the role of electronic phase transitions in enabling non-volatile conductance states.[Vaskivskyi2016_NatCommun] These insights motivate a deeper investigation of switching mechanisms in charge-ordered materials such as $\beta$-Na0.33V2O5.
@@
Upon cooling from room temperature, $\beta$-Na0.33V2O5 undergoes a sequence of phase transitions. Initially, the Na-ions, which occupy one of two distinct crystallographic positions within the tunnel network, [Markina2004] begin to structurally order at $T_{\text{IO}}\sim 230$ K, [Yamada1999, Markina2004] forming a zig-zag pattern that alternates between $\beta$-sites along the $b$-axis. As the temperature is further reduced below $T_{\text{IO}}$, the Na-ion ordering becomes complete, eventually leading to charge redistribution primarily onto the V1 zig-zag chains.[Presura2003, Natanzon2020] Upon further cooling, a charge ordering (CO) transition occurs at $T_{\text{CO}}\sim 130$ K, [van2002, Yamada1999] which is accompanied by the localization of carriers. Optical spectroscopy suggests that these charge carriers may form small polarons, underscoring the significant role of electron–phonon interactions in stabilizing the CO ground state.[Kuntscher2005] At $T_{\text{N}}\sim 22$ K, the system undergoes spin ordering, forming a canted antiferromagnetic structure.[Yamauchi2002] Finally, under hydrostatic pressures exceeding 8 GPa, a superconducting phase emerges below 10 K.[Okazaki2004, Yamauchi2002, Itoh2006, Sirbu2006] The coexistence of strong correlations, reduced dimensionality, and tunable ground states render this system a fertile ground for fundamental studies and for exploring its potential as a platform for memory and computing applications, motivating further investigation of its complex phase behavior.
@@ II Experimental Methods
Single crystals of $\beta$-Na0.33V2O5 were prepared by a solid-state powder synthesis and subsequent melt-growth, as described in .[Parize1989] Na2C2O4 (Sigma-Aldrich) and V2O5 (Sigma-Aldrich) at a 0.18:1 molar ratio (a small stoichiometric excess of Na was used to maximize Na-site filling) were ground together thoroughly in a mortar and pestle and loaded into an alumina combustion boat. The mixture was placed in an ambient-temperature tube furnace, held under vacuum, and heated to 550°C for 16 h. After cooling, the sample was removed, ground, and heated as previously. The identity of the resulting blue-black powder was confirmed by powder X-ray diffraction. To prepare single crystals, the powder product was sealed in an evacuated amorphous silica ampoule, melted at 810°C in a muffle furnace, cooled to 550°C at 2°C/h to crystallize, and cooled to ambient temperature under furnace momentum. Several regular, lustrous black acicular crystals were harvested for measurement.
@@ III RESULTS AND DISCUSSION
Above $T_{CO}$, the data were fit to the nearest-neighbor hopping (NNH) model for small polarons by plotting $\ln(\rho/T)$ versus $T^{-1}$, as shown in Fig. 2(b), following Eq. 1.[Natanzon2020] The extracted activation energy is approximately 63 meV, in agreement with previously reported values for $\beta$-Na0.33V2O5.[Shukla2022, Sirbu2006] This energy represents the thermal barrier for polaron hopping and reflects the degree of carrier localization, making it a critical parameter for understanding the transport mechanism in the high-temperature regime.
@@
However, as shown in Fig. 2(b), the NNH fit is valid only in the temperature range from 130 K to 220 K, suggesting the presence of another transition near this range, which is discussed later. The variable-range hopping (VRH) conduction mechanism below $T_{CO}$ is analyzed by plotting $\ln(\rho/T^{1/2})$ against $T^{-1/4}$, following the temperature dependence described by Eq. 2, [Shukla2022, RN94] as depicted in Fig. 2(c).
@@
This analysis yields Mott’s characteristic temperature $T_{0}$, which is found to be $1.88\times 10^{8}\,\text{K}$. Based on this value, the density of states at the Fermi level $N(E_{f})$, the mean hopping distance $R$, and the mean hopping energy $W$ were calculated. Assuming that the small polaron is localized to a single vanadium site, the localization length was approximated as $\alpha^{-1}\sim 1\,\text{Å}$.[RN94] These calculations yielded $N(E_{f})=1.16\times 10^{21}\,\text{eV}^{-1}\text{cm}^{-3}$, $R=1.27\,\text{nm}$, and $W=100.47\,\text{meV}$. Additionally, for the VRH model to remain valid, the conditions $\alpha R>1$ and $W>k_{B}T$ were verified for all temperatures below $T_{CO}$, as outlined in Eq. 3.[Shukla2022, DKPAUL1973]
@@
To probe the carrier dynamics across the transition, low-frequency noise measurements were performed over the temperature range of 100 K to 300 K. Fig. 2(d) displays representative residual current fluctuations measured above, near, and below the transition, with notably higher fluctuations observed in the vicinity of $T_{\text{CO}}$. The PSD of these fluctuations are shown in Fig. 2(e), exhibiting a clear inverse frequency dependence. Fig. 2(a) shows the PSD magnitude at 0.1 Hz as a function of T, revealing a weak temperature dependence above $T_{\text{CO}}$ and a sharp suppression of noise as the system enters the charge-ordered state below $T_{\text{CO}}$. This weak temperature dependence at high-temperatures is characteristic of hopping conduction in systems [Shklovskii2003, RN97] where transport is dominated by thermally activated hopping between nearest-neighbor localized states. In this regime, the resistivity becomes nearly temperature-independent, and the conduction pathways stabilize, leading to minimal variation in noise magnitude. The strong suppression of noise below $T_{\text{CO}}$, in contrast, reflects the onset of enhanced carrier localization and collective electronic correlations-providing noise-based evidence for the transport crossover from NNH to VRH, as suggested by the resistivity analysis.
Similar to optical measurements showing a gradual ordering of the Na-ions, [Presura2003] further insights into the temperature dependence of the IO and CO can be probed through X-ray diffraction intensities, as shown in Fig. 3 (c). These intensities reveal a gradual increase in CO due to enhanced electron localization as the system is further cooled below $T_{CO}$. This progressive charge localization at lower temperatures underpins for the observed pressure-temperature (PTP-T) diagrams in many other organic and heavy-fermion systems exhibiting pressure-induced superconductivity.
Recent reports have also discussed an additional transition corresponding to Na-ion ordering (IO) at $T_{IO}\sim 240\,\text{K}$.[Obermeier2002, Heinrich2004] Optical measurements have shown that this transition occurs gradually from $T_{IO}$ to $T_{CO}$, where charge localization occurs along the zig-zag chains of vanadium atoms spanning the $b$-axis, centered on the $V_{1}$ vanadium atom.[Yamada1999] Single-crystal X-ray diffraction measurements reveal the atomic structure modulations exhibited by the low-temperature insulating state. $\beta$-Na0.33V2O5 exhibits the C2/m space group at room temperature, shown in Fig. 3 (c). (V1)O6 and (V2)O6 octahedra form V4O16 clusters, which edge-share in chains along $b$, connected along $a$ by shared V1-O2-V1 corners and bridged along $c$ by zig-zag chains of edge-shared (V3)O5 square pyramids. Na-ions occupy paired sites within tunnel-shaped interstices extending along $b$. The proximity of these site pairs (1.953 Å) limits them to half-occupancy, corresponding to a nominal Na1/3V2O5 composition.
@@
Below $T_{\text{CO}}=125$ K, the material adopts an ($a$, $6b$, $c$) supercell, as indicated by the presence of $\mathbf{q}=(h,k/6,l)$ satellite reflections at 100 K in Fig. 3 (b). Vanadium bond valence sum [Bard1999] calculations suggest the formation of an approximate $3b$ periodic CO in conjunction with the extant $2b$ Na ordering as the source of the modulation ( Fig. 3 (a) ). CO has previously been invoked to describe the conductivity and structural transitions and others coincident at 125 K, [van2002, Presura2003] and several charge distribution models have been proposed.[Sirbu2006, Yamaura2002, Itoh2006] Notably, the charge distribution present is nearly identical to that supporting the antiferromagnetic ground state, as determined by single-crystal neutron diffraction below the Néel temperature.[Nagai2005] This magnetic ordering likely coalesces from the CO pattern established at the onset of the semiconductor-insulator transition as presented above. Like the lower-order satellites, Fig. 3 (c) shows that the reflections at $\mathbf{q}=(h,k/6,l)$ intensify with cooling, indicating deepening of charge localization.
@@
In non-interacting systems, noise increases as the temperature decreases due to the narrowing of the density of states near the Fermi energy ($E_{f}$) at lower temperatures, which elevates the energy barriers for hopping. As a result, hopping events between localized states require longer timescales, leading to higher noise magnitudes. Essentially, the reduced availability of states for particles to hop causes higher noise levels at lower temperatures, characteristic of non-interacting systems.[RN97, Burin2006] In contrast to characteristic non-interacting patterns, the observed noise behavior in single crystals of $\beta$-NaxV2O5 exhibits pronounced differences. Indeed, the noise magnitude decreases below the characteristic transition temperature TCO, a phenomenon that cannot be explained within the framework of non-interacting models. The observed anomaly is attributed to the onset of strong electron-electron correlations due to CO, which drive a collective reorganization of localized states under the influence of Coulomb interactions. In this correlated regime, these interactions suppress the independent hopping fluctuations characteristic of non-interacting systems, ultimately giving rise to a glassy state.[Shtengel2003, RN100] As previously reported in optical studies, the observed mid-infrared absorption, combined with strong electron-phonon coupling and effective carrier density, suggests that small polaron transport plays a dominant role in $\beta$-Na0.33V2O5.[Presura2003] Hall measurements have revealed that the total number of charge carriers remains largely unaffected by temperature changes in the range of 189 K to 298 K, with the carrier density approximately matching the concentration of Na atoms in the crystal.[Perlstein1968] This gives additional support to the polaronic nature of conduction, as each excess electron introduced into the lattice by the intercalated monovalent Na localizes on the $V_{1}/V_{2}$ vanadium sites, which exhibit a mixed valence of +4/+5.
@@
Further analysis is presented in Fig. 5(c) and (d). In Fig. 5(c), the noise magnitude at $f=10~\mathrm{mHz}$ remains low at low voltages, but increases sharply as the system approaches $V\sim 16.5~\mathrm{V}$, peaking just before the current discontinuity. The abrupt rise abrupt rise in noise is indicative of critical dynamics often observed during the collapse of an ordered electronic state. In the context of $\beta$-Na0.33V2O5, such dynamics are consistent with the field-driven destabilization of the charge-ordered state formed by localized small polarons. It suggests a regime where the applied voltage competes with the intrinsic pinning potential imposed by lattice or impurity interactions. Simultaneously, Fig. 5(d) shows the evolution of the PSD exponent $\alpha$, which is close to 1 at low voltages-typical of $1/f$ noise. As the transition is approached, $\alpha$ increases toward 2, reflecting a shift to noise behavior dominated by slow, diffusive-like dynamics. A $1/f^{2}$ noise spectrum ($\alpha\approx 2$) is indicative of non-equilibrium dynamics associated with depinning or cooperative motion. Similar behavior has been observed in charge-ordered organic conductors such as $\kappa$-(BETS)2Mn[N(CN)2]3, where a sharp rise in noise and $\alpha\approx 2$ appears below the transition, reflecting slow, correlated fluctuations of polar nanoregions.[Thomas2024]
@@
Moreover, the temperature dependence of the threshold voltage $V_{\text{th}}$ shown in Fig. 6(b), displays an exponential behavior. This trend is consistent with prior observations in related charge-ordered systems,[Yamanouchi1999, Asamitsu1997] where the CO phase becomes increasingly susceptible to electric-field-induced breakdown at lower temperatures.